One number is fifteen more than twice another. If their sum is decreased by nine, the result is eighteen. Find the numbers. (Enter solutions from smallest to largest.)
Answer by ewatrrr(24785)


Hi
One number is fifteen more than twice another
Let x and (2x+15) represent the two numbers
Question states***
x + (2x+15)-9 = 18
Solving for x
3x = 12
x = 4
(2x+15)= 23 The two numbers are 4 and 23.
CHECKING our Answer***
(4+23) - 9 = 27 - 9 = 18